OpenStudy (anonymous):

Simplify following expression : -3r+7-3r-12

jimthompson5910 (jim_thompson5910):

Were you able to get anywhere with this?

OpenStudy (anonymous):

Sort of .. It was one of my classwork problems and I got it wrong .

jimthompson5910 (jim_thompson5910):

what did you get?

OpenStudy (anonymous):

I dont have my paper , just the equation . So let me solve it again and see ,

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

i got r-19 , this was a re-due , so im not sure if that was my very first answer but just resolving it I got : r-19

jimthompson5910 (jim_thompson5910):

what is -3r - 3r

OpenStudy (anonymous):

0 ?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

so if it's 0 , wouldn't it just be ' r ' ?

jimthompson5910 (jim_thompson5910):

you're thinking of 3 - 3

jimthompson5910 (jim_thompson5910):

or -3 + 3

jimthompson5910 (jim_thompson5910):

-3 - 3 = -6

jimthompson5910 (jim_thompson5910):

so that's why -3r - 3r = -6r

jimthompson5910 (jim_thompson5910):

7 - 12 = -5, so put together, we can say all this -3r+7-3r-12 -3r - 3r + 7 - 12 (-3r - 3r) + (7 - 12) (-6r) + (-5) -6r - 5

OpenStudy (anonymous):

so -3r+7-3r-12 simplified would be -6r-19 ?

jimthompson5910 (jim_thompson5910):

no, it would be -19 IF it was -7 - 12 but it's not

OpenStudy (anonymous):

i mean -6r-5 ...

jimthompson5910 (jim_thompson5910):

it's +7 - 12 = 7 - 12 = -5

jimthompson5910 (jim_thompson5910):

better

OpenStudy (anonymous):

thank you so much:)

jimthompson5910 (jim_thompson5910):

sure thing